Please use excel and state the formula used and what was inputted to solve it, thank you

1. A budgeting Web site reported that 20 % of U.S. households have withdrawn money from a 401(k) or other retirement account for needs other than retirement in 2013. A random sample of 11 U.S. households was selected.

What is the probability that exactly three households withdrew funds from a retirement account for needs other than retirement?

Find the mean, variance, and standard deviation of the binomial distribution with the given values of n and p.

n=60

p= 0.8

A government's department of transportation reported that in 2009, airline A led all domestic airlines in on-time arrivals for domestic flights, with a rate of 83 %.

What is the probability that in the next six flights, exactly four flights will be on time?
Answer #1

Hello Sir/ Mam

We are most delighted to answer, but I would request you to please ask atmost 1 question or 4 subparts per post. Your cooperation will be highly appreciated. Thanks!

Q - 1 -

Given that 20% of the households have withdrawn funds from a retirement account for needs other than retirement.

Hence, the probability that the randomly selected household has withdrawn funds from a retirement account for needs other than retirement = 0.2

Now, we have to find  the probability that exactly three households withdrew funds from a retirement account for needs other than retirement.

It is a simple case of Binomial Distribution in which n = 11, p = 0.2, x = 3 and hence we can apply formula :

(:)(p)(I-P)-r-(,) (0.2)3(0.8)8 = 22.146%

To do so in excel, use the following formula, " =COMBIN(11,3)*(0.2^3)*(0.8^8)" and answer comes out to be 22.146%

2.

n = 60

p = 0.8

q = 1 - p = 0.2

Mean = n*p = 60 * 0.8 = 48

Variance = n*p*r = 60 * 0.8*0.2 = 9.6

Standard \ Deviation = \sqrt{n*p*r} = \sqrt{60 * 0.8*0.2 }= \sqrt{9.6}=3.0984

I hope this solves your doubt.

Feel free to comment if you still have any query.

Do give a thumbs up if you find this helpful.


answered by: ANURANJAN SARSAM